GATEWAY MEDIATED INTERSYSTEM CROSSING IN C_{2}H_{2}

\underline{KEVIN L. CUNNINGHAM}, STEPHEN DRUCKER, SELEN ALTUNATA, CHRISTOPHER G. MORGAN, and ROBERT W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge MA 02139-4307.

Recent Surface Electron Ejection by Laser-Excited Metastable (SEELEM) spectra have provided a clearer picture of the intersystem crossing of the 3ν_{3} level of the S_{1} surface of C_{2}H_{2} to near-degenerate triplet states. The simultaneously recorded LIF and SEELEM spectra indicate that a single vibrational level of the T_{3} surface couples the 3ν_{3} level of the S_{1} surface to a dense manifold of T_{1} and T_{2} states. This proposed coupling mechanism, Gateway Mediated Intersystem Crossing (GMISC), and other proposals will be discussed in light of the new data.